Determine the temperature distribution u(x, t) in a bar of length L, whose ends are kept at zero

temperature, and whose initial temperature is:

i) u(x, 0) = f(x) = U0

ii) u(x, 0) = f(x) = k sin πx/L



1
Expert's answer
2022-01-10T17:08:08-0500

We shall apply the heat equation. Which is:

"U_t=cU_{xx}"

To solve this, let "U=XT\implies U_t=XT',U_{xx}=X''T"

The equation becomes:

"XT'=cX''T\\
\frac{T'}{cT}=\frac{X''}{X}=-\lambda^2"

Solving the two differential equations, we have that:

"T=Ae^{-c\lambda^2t} \text{ and } X=B\cos \lambda x+ C \sin \lambda x"

Thus, "U(x,t)=XT'=(A_1\cos \lambda x+ A_2 \sin \lambda x)e^{-c\lambda^2t}"

Applying the boundary conditions "U(0,t)=U(L,t)=0"

"U(0,t)= A_1e^{-c\lambda^2t}=0\\
\text{But, exponential does not vanishes}\\
\implies A_1=0\\
U(x,t)=A_2e^{-c\lambda^2t}\sin \lambda x"

Applying the second condition;

"U(L,t)=A_2e^{-c\lambda^2t} \sin \lambda L=0\\
\implies A_2\sin \lambda L=0\\
\text{But, } A_2≠0\\
\implies \sin \lambda L=0\implies \lambda=\frac{n \pi}{L}, n\in \Z\\
\text{Thus, } U(x,t)=A_2e^{-c(\frac{n \pi}{L})^2t}\sin \frac{n \pi}{L}x~~~~~~~~~~~~~~~~(*)"

Now to apply the conditions given in the question.

(i) "U(x,0)=f(x)=U_0"

By superposition principle, (*) becomes

"U(x,t)=\u2211_{n=0}^\\infty\u200bA_n\u200be^{\u2212c(\\frac{n \\pi}{L})^2t}\\sin \\frac{n \\pi}{L}\u200bx\\\\\nU(x,0)=\\sum_{n=0}^\\infty A_n\\sin \\frac{n \\pi}{L} x=U_0"

This is nothing but the Fourier sine series of "U_0"

Where,. "A_n=\frac{2}{L}\int_0^LU_0 \sin\frac{n \pi}{L}x ~dx=\frac{2}{n\pi}U_0[1-(-1)^n]"

Thus "U(x,t)=\sum_{n=0}^\infty \frac{2}{n\pi}U_0[1-(-1)^n] e^{-c(\frac{n \pi}{L})^2t}\sin \frac{n \pi}{L} x\\"


(ii) For the condition "U(x,0)=f(x)=k \sin \frac{ \pi}{L}x"

"U(x,0) \text{ from } (*) is\\
U(x,0)=A_2\sin \frac{n \pi}{L}x= k \sin \frac{\pi}{L}x\\
\text{Comparing both sides}\\
A_2=k, n=1\\
\text{Thus, } U(x,t)=ke^{-c(\frac{\pi}{L})^2}\sin \frac{\pi}{L}x"